. In 1985, 45 percent of a document storage facility’s 60 customers were banks, and in 1987. 25 percent of its 144 customers were banks. What was the percent increase from 1985 to 1987 in the number of bank customers the facility had?
(A) 10.7%
(B) 20%
(C) 25%
(D) 33.3%
(E) 58.3

Answer by longjonsilver(2297)

find how many banks there were in 1985 and 1987. These 2 numbers would increase. We then find the % increase between these two.

So, in 1985: .45*60 = 27 banks
in 1987, .25*144 = 36 banks

increase is 9 banks, from the original of 27, so % increase = (9/27)*100 which is 33.3333333%: answer D.

jon.